Given: C = 100 + 0.75Yd (where Yd = Y-T) I = 120-600i G = 200 T = 20 + 0.2Y Ms/P = 300 Md/P = 50+0.5Y-600i 


1
Expert's answer
2021-05-17T09:57:54-0400

"Y = C + I + G = 100 + 0.75(Y - 20 - 0.2Y) + 120 - 600i + 200 = 0.6Y + 405 - 600i."

0.4Y = 405 - 600i,

Y = 1,012.5 - 1,500i.

Ms/P = Md/P,

50 + 0.5Y - 600i = 300,

0.5Y = 250 + 600i,

Y = 500 + 1,200i.

1,012.5 - 1,500i = 500 + 1,200i,

2,700i = 512.5,

i = 0.1898 or 18.98%.

Y = 500 + 1,200×18.98 = 727.78.